OpenStudy (anonymous):

The lengths of the sides of a rectangle become 4 times greater than the original measurements. Which statement is true about the rectangle?

OpenStudy (anonymous):

Its area becomes 4 times greater, and perimeter becomes 20 times greater. Its area becomes 16 times greater, and perimeter becomes 4 times greater. Its area becomes 4 times greater, and perimeter becomes 16 times greater. Its area becomes 20 times greater, and perimeter becomes 4 times greater

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let L be the length and W be the width The area is then A = LW and the perimeter is P = 2L + 2W

OpenStudy (anonymous):

so what do i plug in ?

jimthompson5910 (jim_thompson5910):

If you multiply each side by 4, then L becomes 4L and W becomes 4W

jimthompson5910 (jim_thompson5910):

So the new area is then A = LW A = (4L)*(4W) ...I'll let you continue

OpenStudy (anonymous):

is it the third one?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@huneya am I right?

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

........uh

jimthompson5910 (jim_thompson5910):

A = LW A = (4L)*(4W) A = (4*4)*(LW) A = 16LW So the new area is 16 times greater

OpenStudy (anonymous):

so I'm right